find the slope of the line parallel to the line 4x-2y-1
Answer by galileo(7)

The slope-intercept formula of a line is y = mx + b where m is the slope and b is
the y-intercept
4x-2y-1 = 0
-2y = -4x + 1
y = 2x - 1/2
The slope of the line is 2 and a y-intercept of -1/2. Any line parallel to this line must also have a slope of 2.